OpenStudy (anonymous):

test the series for convergence or divergence. Use comparison or limit comparison test.

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} \frac{ \ln n }{ n }\]

OpenStudy (anonymous):

@jim_thompson5910 could you help me please

OpenStudy (anonymous):

i got an to be \[\frac{ \ln n }{ n } and.....bn=\frac{ 1 }{ n}\]

OpenStudy (anonymous):

i am not sure if i am doing it right or not

OpenStudy (anonymous):

I suggest the integral test

OpenStudy (anonymous):

i cant do this using the comparison or limit test?

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (anonymous):

alright

OpenStudy (dan815):

ok u know that 1/n^p p has to be atleast greater than 1 to converge

OpenStudy (anonymous):

yeah

OpenStudy (dan815):

and ln n / n > 1/n

OpenStudy (anonymous):

yes, you can use limit comparison. compare 1/n to ln(n)/n \[\lim_{n \rightarrow infinity } (b _{n}/a _{n})\] bn is 1/n an is ln n/n 1/n * n/ln(n) = lim 1/ ln(n) = infinity, not defined. So diverge.

jimthompson5910 (jim_thompson5910):

You have the correct \(\Large a_n\) and \(\Large b_n\) Solve ln(n)/n = 1/n, to get ln(n)/n = 1/n ln(n) = 1 n = e^1 n = e n = 2.718 (approximately) Notice how if n > 2, say n = 3, then we will have ln(n)/n = ln(3)/3 = 0.3662 1/n = 1/3 = 0.33333 So this shows us that 1/n < ln(n)/n for all n such that n > 2 You can check other values of n that are larger than 2 and you'll find 1/n < ln(n)/n holds true

OpenStudy (rational):

limit comparison test is nice here as it avoids all above algebra simply divide and take the limit

jimthompson5910 (jim_thompson5910):

you can use a p-series test to find that 1/n diverges Let \[\Large a_{n} = \frac{\ln(n)}{n}\] \[\Large b_{n} = \frac{1}{n}\] Since \(\Large b_{n} < a_{n} \) for n > 2 (n is an integer), and the b sequence diverges, this means the 'a' sequence must also diverge as well

jimthompson5910 (jim_thompson5910):

Either method works, so you can take your pick.

OpenStudy (anonymous):

so what if its \[\frac{ 1 }{1+e^n }\]

OpenStudy (anonymous):

would bn still equal 1/n?

OpenStudy (rational):

whats the full question ? 1/(1+e^n) is not a series

OpenStudy (anonymous):

its \[\sum_{n=1}^{\infty} \frac{ 1 }{ 1+e^n }\]

OpenStudy (anonymous):

i was thinking maybe bn = \[\frac{ 1 }{ e^n } \]

OpenStudy (rational):

try comparing 1/(1+e^n) with 1/n^2

OpenStudy (rational):

depends on what series you already know converges

OpenStudy (rational):

if you already know that SUM 1/e^n converges then you may use that for comparison

OpenStudy (rational):

i prefer 1/n^2 though as it is more popular

jimthompson5910 (jim_thompson5910):

You can also use the integral test

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+e^n } \ge \frac{ 1 }{ n^2 }\]

OpenStudy (rational):

ou're doing it reverse For convergence using comparison, you need to show that the terms in sequence {1/(1+a^n)} are LESS than the terms in sequence {1/n^2}

OpenStudy (anonymous):

i can but my teacher wants us using the comparison or limit comparison test

OpenStudy (rational):

for \(n\gt 0\) we have \[e^n+1 \gt n^2 \implies \dfrac{1}{e^n+1} \lt \dfrac{1}{n^2} \] since the series \(\sum \frac{1}{n^2}\) converges, the series \(\sum \frac{1}{e^n+1}\) also converges by direct comparison test.

OpenStudy (anonymous):

how do you use the direct comparison test in this problem?

OpenStudy (rational):

"direct comparison test" is another name for "comparison test"

OpenStudy (rational):

both are same

OpenStudy (anonymous):

oh okay and when do you know its divergence?

OpenStudy (rational):

lets see a quick example : 64 + 32 + 16 + 8 + 4 + 2 + ... what do you know about convergence/divergence of above series ?

OpenStudy (anonymous):

the series is decreasing

OpenStudy (rational):

right, thats a good reason to suspect that the series can possibly converge

OpenStudy (rational):

careful when you say : "the series is decreasing" that statement makes no sense

OpenStudy (rational):

i know you meant "the terms of series are decreasing"

OpenStudy (anonymous):

like for example this one \[\sum_{n=1}^{\infty} \frac{ 1 }{ n^n}..... diverges\]

OpenStudy (anonymous):

because \[\frac{ 1 }{ n^n } > \frac{ 1 }{ n^2 }\]

OpenStudy (anonymous):

right?

OpenStudy (rational):

\[n^n \gt n^2 \implies \frac{1}{n^n} \lt \frac{1}{n^2}\] so the series SUM 1/n^n actually converges

OpenStudy (rational):

10 > 2 so 1/10 < 1/2

OpenStudy (rational):

2 > 1 so 1/2 < 1/1

OpenStudy (anonymous):

but look if n =3 3^3>3^2

OpenStudy (anonymous):

right?

OpenStudy (rational):

right, 3^3 > 3^2 so we have 1/3^3 < 1/3^2

OpenStudy (anonymous):

oh yeah forgot they were fractions

OpenStudy (anonymous):

so no need to find the limit for this one since we already know it converges right?

OpenStudy (rational):

comparison test doesn't require working the limit can we go back to earlier example series and finish it off ?

OpenStudy (anonymous):

oh yes

OpenStudy (rational):

``` 64 + 32 + 16 + 8 + 4 + 2 + ... ``` what do you know about convergence/divergence of above series ?

OpenStudy (anonymous):

well the term of the sequence is decreasing by 32/n or something like that

OpenStudy (rational):

right so what can you say about its convergence ?

OpenStudy (anonymous):

the sequence and series converges?

OpenStudy (rational):

why ?

OpenStudy (anonymous):

i dont really know

OpenStudy (rational):

not covered geometric series yet ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but could you tell me what you mean?

OpenStudy (rational):

``` 64 + 32 + 16 + 8 + 4 + 2 + ... ``` suppose we know that that series converges then the below series also converges by "comparison test" because each term in below series is positive and is less than the corresponding term in above series which converges. ``` 63 + 31 + 15 + 7 + 3 + 1 + ... ```

OpenStudy (rational):

i was just trying to show you why "comparison test" works but that example wont make sense if "geometric series" is not studied yet

OpenStudy (anonymous):

is geometric series similar to comparison test?

OpenStudy (rational):

geometric series is name of an actual series comparison test is a test for knowing whether a series converges or diverges both are different things

OpenStudy (anonymous):

could you show me an example of a geometric series?